4x + 10y = -52
-3y =14x -9
12y - 2x = 17
2x = 9y + 4
8x +13y = 49
Tell me what all of these are in Standard form please?

Answer by ewatrrr(24785)

 
Hi
Using the standard form of a line as in the form Ax + By = C where A is >0
4x + 10y = -52 OK
-3y =14x -9 14x + 3y = 9
12y - 2x = 17 2x - 12y = -17
2x = 9y + 4 2x -9y = 4
8x +13y = 49 OK